If 6 is added to a number and than subtracted from the same number, the product of the sum and the difference is 189 . What is the number?
Answer by stanbon(48575)

If 6 is added to a number and than subtracted from the same number, the product of the sum and the difference is 189
------------------------
Let the number be "x".
EQUATION:
(x+6)(x-6) = 189
x^2 -36 = 189
x^2 = 225
x = 15 or x = -15
======================
Cheers,
Stan H.